# GUI-Bee [Uncaptioned image]     : Align GUI Action Grounding to Novel Environments via Autonomous Exploration

### v1

@@ Appendix D Predicting $\hat{Q}(a^{t}_{x})$ with MLLM through In-Context Learning
To predict $\hat{Q}(a^{t}_{x})$ for candidate actions $a^{t}_{x}$, in the text prompt, we include the MLLM with the natural language escriptions $D^{t}$ approximating current state $S^{t}$, two example actions $(a_{\text{eg1}},a_{\text{eg2}})$ and their Q-values $(Q(a_{\text{eg1}}),Q(a_{\text{eg2}}))$, along with the visual input using the Set-of-Mark method (Yang et al., 2023). Figure 10 shows an example of the full input and output.
@@ Appendix E Experiment Details for Exploration and Fine-tuning
We fine-tune three GUI grounding models—SeeClick (Cheng et al., 2024), Qwen-GUI (Chen et al., 2024), and UIX-7B (Liu et al., 2024)—using the data generated by the GUI-Bee agent. Fine-tuning is performed in two input configurations: vision-only, where the input consists of GUI screenshots only, and Vision+A11y, where the input includes both GUI screenshots and the accessibility tree embedded in the text prompt.
@@
SeeClick and Qwen-GUI are based on Qwen-VL (Bai et al., 2023), while UIX-7B is derived from Llava-1.6 (Li et al., 2024) with Qwen2-7B-Instruct (Yang et al., 2024) as the primary LLM backbone. For models that predict bounding boxes, such as Qwen-GUI and UIX-7B, the center of the predicted bounding box is used as the final output point for evaluation.
@@ Appendix G Examples of NovelScreenSpot
We randomly sample data from each environment in the NovelScreenSpot benchmark and present examples in Figure 11, Figure 12, Figure 13, Figure 14, and Figure 15. Each figure illustrates the GUI screen, the A11y string, the corresponding query, and the ground truth target element, showcasing the diversity and environment-specific nature of the benchmark.

### v2

@@ 6 Conclusion
7 Limitations
The GUI-Bee agent excels in tailoring GUI action grounding models for specific environments but has limitations. It uses Multimodal Large Language Models (MLLMs) like GPT-4o, which provide high-quality data but come with higher computational costs, latency, and privacy issues. Additionally, as exploration within a specific environment continues and the number of exploration steps increases, the GUI-Bee agent accumulates an increasingly long history of exploration data. Processing this longer history can introduce additional overhead, increasing the time required to select actions for subsequent exploration steps. This may impact scalability in highly complex environments or during extended exploration sessions.
@@ Appendix D Predicting $\hat{Q}(a^{t}_{x})$ with MLLM through In-Context Learning
To predict $\hat{Q}(a^{t}_{x})$ for candidate actions $a^{t}_{x}$, in the text prompt, we include the MLLM with the natural language escriptions $D^{t}$ approximating current state $S^{t}$, two example actions $(a_{\text{eg1}},a_{\text{eg2}})$ and their Q-values $(Q(a_{\text{eg1}}),Q(a_{\text{eg2}}))$, along with the visual input using the Set-of-Mark method Yang et al. (2023). Figure 10 shows an example of the full input and output.
@@ Appendix E Experiment Details for Exploration and Fine-tuning
We fine-tune three GUI grounding models—SeeClick Cheng et al. (2024), Qwen-GUI Chen et al. (2024), and UIX-7B Liu et al. (2024)—using the data generated by the GUI-Bee agent. Fine-tuning is performed in two input configurations: vision-only, where the input consists of GUI screenshots only, and Vision+A11y, where the input includes both GUI screenshots and the accessibility tree embedded in the text prompt.
@@
SeeClick and Qwen-GUI are based on Qwen-VL Bai et al. (2023), while UIX-7B is derived from Llava-1.6 Li et al. (2024) with Qwen2-7B-Instruct Yang et al. (2024) as the primary LLM backbone. For models that predict bounding boxes, such as Qwen-GUI and UIX-7B, the center of the predicted bounding box is used as the final output point for evaluation.
@@ Appendix G Examples of NovelScreenSpot
We randomly sample data from each environment in the NovelScreenSpot benchmark and present examples in Figure 11, 12, 13, 14, and 15. Each figure illustrates the GUI screen, the A11y string, the corresponding query, and the ground truth target element, showcasing the diversity and environment-specific nature of the benchmark.